OpenStudy (anonymous):

I am trying to find the taylor series for (x,y) = (x+y)^n. I began by getting the partial derivatives as df/dx = n(x+y)^(n-1) (1+y) and df/dy=n(x+y)^(n-1) (x+1). don't know where to go from there.

OpenStudy (amistre64):

yeah, the taylor of multiple variables takes partials and attributes then to powers ....

OpenStudy (amistre64):

and it might be easier to keep track of if you notate it as: Fx, Fy, Fxx, Fyy, Fxy, Fyx, ...

OpenStudy (anonymous):

sorry, need a little more guidance as to where to go from here.

OpenStudy (amistre64):

well, from what i recall, the subscripts help to define the variable exponents Fx is the coeff of x Fxx is the coeff of x^2 Fxy is the coeff of xy etc

OpenStudy (amistre64):

in a single variable setup that relates to Fx is coeff of x Fxx is coeff of x^2 Fxxx is coeff of x^3 etc ....

OpenStudy (amistre64):

(x+y)^n \[ F_x = n(\hat x+\hat y)^{n-1}\\ F_{x^{(2)}} = n(n-1)(\hat x+\hat y)^{n-2}\\ F_{x^{(3)}} = n(n-1)(n-2)(\hat x+\hat y)^{n-3}\\ F_{x^{(k)}} = n(n-1)(n-2)...(n-k+1)(\hat x+\hat y)^{n-k}\\ \] the Fys will have the same structure of course its the intermediarys of xy x^2y xy^2 and such that can get harrowing